If the point P is on the graph of a function f, find the corresponding point on the graph of the...

P(3, -2); y = 2f(x - 2) + 1
Answer #1

Point on y= f(x) : P(3, –2)
Transformed function: y = g(x) = 2f(x – 2) + 1

Transformations (that transform f into g):
• a vertical stretch by a factor of 2: ... P(3, –2) ➞ P⸍(3, –4)
• a horizontal slide right 2: ... P⸍(3, –4) ➞ P⸍⸍(5, –4)
• a vertical slide up 1: ... P⸍⸍(5, –4) ➞ P⸍⸍⸍(5, -3)

∴ Corresponding point on new graph is P⸍⸍⸍(5, –3)

answered by: Smiley (I NEED MS.SUE OR WREITEACHER!!) HELP!!
Answer #2
Point on y= f(x) : P(3, –2)
Transformed function: y = g(x) = 2f(x – 2) + 1

Transformations (that transform f into g):
• a vertical stretch by a factor of 2: ... P(3, –2) ? P?(3, –4)
• a horizontal slide right 2: ... P?(3, –4) ? P??(5, –4)
• a vertical slide up 1: ... P??(5, –4) ? P???(5, –3)

? Corresponding point on new graph is P???(5, –3)
answered by: inters